SOHO SPWG Minutes - April 1996

April 19, 1996, SOHO EAF, GSFC
Chair: Piet Martens
Notes: Laura Allen and Eliane Larduinat


V. Domingo, B. Fleck, A. Gabriel, K. Bocchialini, M. Huber, J.-P. Delaboudiniere, B. Schmieder, P Lemaire, D. Spadaro, J. Raymond, K. Dere, B. Kliem, B. Thompson, J. Newmark, J. Saba, R. Kavanagh, J. Behuncik, A. Galvin, O. Kjeldseth-Moe, R.D. Bentley, D. Muhonen, L. Allen, C. Froehlich, L. Sanchez, E. Larduinat, A. Poland, P. Martens, J.-L. Bertaux.


  1. Review of Action Items
  2. Status update
  3. The coming month:
  4. JOPs status and priorities
  5. New JOPs
  6. Experiment anomaly reporting
  7. Publication of SOHO schedules
  8. Summary data
  9. As-run data-base
  10. Status of data distribution
  11. Special issue: First results from SOHO
  12. AOB

1. Actions

SPWG 4.1:
Piet Martens to discuss duty cycle adjustments with Maryse Hervieux, and report back on future plans to SPWG.

SPWG 4.2:
On Piet Martens to finalize and coordinate roll maneuver timeline with FOT.

SPWG 4.3:
On the Project Scientist to request officially from MDI PI to perform jitter tests for each momentum management as Matra has requested this information to do trending on the reaction wheels

SPWG 4.4:
Luis Sanchez to hold meeting with instrument reps to discuss as-run database, summary data, and populating archive

1a. Reviewing action items from last meeting

SPWG 3.1:

SPWG 3.2:
On-going, now SPWG 4.2

SPWG 3.3:

SPWG 3.4:
Plans for ECS to send summary data to CDHF for CD-ROMs. Summary data does not need to be calibrated. Only EIT has given summary data to date. VIRGO is going to check with Antonio for summary data. MDI will supply magnetograms. UVCS will supply by end of April. SUMER and CDS placing rasters -- no concrete plans for data delivery.

SPWG 3.5:
On-going, now SPWG 4.3

2. Status Update

Accumulating long data set for the last 2 - 3 weeks. Been working in a mode (no polarizers) in the blue wing of light. This will enhance the data, there will be no losses due to mechanism problems. Not sure what is wrong concerning the mechanisms; but will continue in this manner for a while

Working -- lots of good results. Want to see how measurements made in different directions (N,S,E,W) compare.

Spacecraft Heater Discussion

At the beginning of the week brought heaters to 50 % . (PTCS Adjustment) VIRGO asked why were the heaters so abruptly changed from 100% to 50% ? Especially when seasonal differences account for only a 3% change... VIRGO thought that 100% was the range of the mission.

From here on, to correct for seasonal changes, duty cycle will be adjusted in 10% steps, if needed, every 8 weeks during maneuvers.

Continuous structure program. Uploading flat field tables by background queue, had some interference with NRT commanding instrument-wise so they stopped. Dynamics campaign will be getting high rate data. Lot of Helioseismology data taken. Showed magnetogram of active region taken on March 13. Showed image taken in continuum. Think of a better way to coordinate with coronal instruments to support them.

Working well. Noticed sensitivity decline of detector to high counts. There was a 30% change in a reference spectra but this is being corrected for. Think this is due to the microchannel plate. May change to other detector in July. However, with this change there will be a gain of 20% in efficiency. SUMER is trying to avoid active regions -- trying to select lines to study that have ~ 10 counts /sec. Permits long deep exposures in the corona at 1.5 Solar Radii

Galactic Map of Lyman-alpha shown; Solar wind higher in Southern than in the Northern Hemisphere; but depends on calibration between the 2 sensors. Overall sky brightness has been increasing since December, however still need to determine if this is real or detector effect. Also problem that Lyman-alpha intensity plagued by Milky Way stars/diffuse light. In Northern Hemisphere have Barium Fluoride window that excludes Lyman-alpha; but there is no pixel like this on the Southern detector. One way to discrimate is if the pattern of stray light is the same after the roll.. Despite Sun shade, see a lot of light.. stray? coronal emission?

SWAN's main objectives during a S/C roll are:

  1. Check cross calibration of the 2 SWAN sensors
  2. Near Sun observations: Discrimate Stray Light from genuine coronal Lyman-alpha emission.
  3. Subtract from Total Emission what has been recorded from Star/Galactic Contribution
  4. Monitoring of Comets Infalling on the star Beta Pictorus Put Barium Fluoride pixel on the star and stay there for some time (~ one event per day)

Thinks instrument is aging. Full sky map takes ~ 22 hours.

There was a lot of discussion about whether or not to rotate the spacecraft as SWAN wants to. EIT had nothing against the rotation. VIRGO and MDI strongly opposed it stating that during the Commissioning phase this would have been okay, but now we are in the observational period and the main goal is to provide CONTINUITY of the observations. VIRGO can't compensate for the rolls. Data of the LOI cannot be used during the roll. Same for MDI. MDI wants only minutes of interruption -- the timing is especially bad with the two months of continuous contact being so close -- MDI has only 3 chances to do this. Coronal instruments would have to close doors. CDS reconfiguration takes 2 days. Art Poland suggested since we have not gotten a smooth flow yet planning these operations we should wait.

Discussion also involves JOP002, which requires + or - 90 degrees roll. Requirement: slit of 2 spectrometers, only a polar hole will have extent and depth to permit this. Desirable to do this soon, closer to solar minimum. Choice of North or South pole based on previous days' observations.

Requirement: turn 90 degrees and stay there for the sequence. at first scheduled for 13 hours, then made to 17 hours. We now want the duration to get back to the original 13 hours, to limit total duration to 16 hours

AGREEMENT: If tests of JOP002 in the upcoming weeks go well, then run JOP002 on May 20, as per the timeline draft of SOC. If tests do not run well, then run JOP002 in July. Total duration, including roll maneuvers not to exceed 16 hours. Exercise may be repeated in September, or later, if needed, but this can be discussed/decided in the upcoming SWT. Piet Martens to finalize timeline with FOT to have them arrange for DSN backups, etc.

Nominal. Campaigns: flows, bursts, jets, XBP Found high velocity events. GIS shows some degradation at 304 line and needed to adjust the voltage. It is not unexpected since it is such a strong line they expected problems.

Nominal. Considers itself still in commissioning phase due to S/W. New OBE load for LASCO software. Still no planning tool yet. EIT is happy with the repointing, no need for extra aluminum filter. Do not plan on using different telemetry submodes soon, their S/W can't handle it yet.

Synoptic observations (taking 55% of the time), running sequences devoted to line profiles, solar wind velocities, optimized sequence of commands for JOP002. No results yet from temperature and line profiles. UVCS sees most of lines they expected to see. UVCS sees more lines than they originally thought.

Waiting for energetic particle event for STOF. Getting data at low levels. Still commissioning, got flight software patches.

Failure 6 weeks ago with first telescope. Second telescope has higher noise. Nominal, operating and taking data now.

3. Coming month

XBP Campaign needed better coordination with DSN schedule In future, asking for 3 weeks ahead notice for planning future campaigns such as XBP.

Request was made to be better informed of project database updates and when they are scheduled.

Plans for 2 months continuous

Information on collaborations with Ground-based observatories, coordination contacts, etc. should be forwarded to SOC for inclusion into the schedule.

MDI will have no high resolution observations, just full disk magnetograms every 96 minutes.

On June 10th, Venus passes in front of the solar corona, not the disk.

MDI jitter test during momentum management. Monitor jitter performance of wheels but have not received a formal request to do this. (action on VD). Found the amplitude of the resonances to be between 0.001 - 0.02 arcseconds. MDI wants this done in very close proximity and is currently performing this before and after the movement (but not during..).

4. JOPs and Priorities

The following JOP schedule was determined at the meeting. See the SOHO Web pages for updates to the schedule! (

For latest updates on JOP descriptions, see also the SOHO Web pages: (

UPDATE: April 22, 1996
                 MAY/JUNE SCHEDULE                              

M Apr 22 (W17)  
T Apr 23        
W Apr 24        START MDI Continuous, JOP002 (test only no rolls -- 3 hours)
T Apr 25        JOP002 (contingency), JOP020: Nanoflares
F Apr 26        JOP028 (4 hours); JOP012 6:00-12:00 UT SUMER Prominences (test)
S Apr 27        Intercal 002, phase 2
S Apr 28        END MDI Continuous

M Apr 29 (W18)  JOP035 (depending on opportunity)
T Apr 30        JOP013 (TBC) LASCO/SWAN start Comet Observation
W May 01        JOP012 Prominences (test) or filaments
T May 02        JOP015 (during NRT) 
F May 03        JOP002 (test; no rolls)
S May 04
S May 05

M May 06 (W19)  LASCO end Comet Observations
T May 07        
W May 08      
T May 09        JOP015 
F May 10        Special ISTP configuration (WIND, Geotail)
S May 11
S May 12

M May 13 (W20)  
T May 14        JOP022 (TBC)
W May 15 
T May 16        JOP027  (TBC)   
F May 17        Momentum management -- some experiments close doors
S May 18        Experiments reopen doors, CDS reconfigure
S May 19        Normal operations, CDS reconfigure

M May 20 (W21)  S/C 90 degrees roll (JOP 002, SWAN galactic map; 16 hours max)
T May 21        Second momentum management (if needed)
                Orbit correction burn
                Third momentum management (if needed)
W May 22        Contingency, experiments reopen doors   
T May 23
F May 24        Start of 2 Months Continuous    
S May 25
S May 26        Star calibration (SUMER/UVCS)

M May 27 (W22)
T May 28        JOP031 (streamers) (TBC.  Could be delayed)
W May 29        JOP031    
T May 30        JOP031
F May 31
S Jun  1
S Jun  2 
Are there any other JOPs in which Yokhoh participation is desired? JOP leaders should coordinate with them. Bob Bentley volunteers for this.

A schedule for Intercalibration JOPs will be determined by the SICWG.

5. New JOPs

JOPS updated on the Web. Intercal2 has been updated. No new JOPs proposed in last week.

Phil Judge and Olav (CDS) are working on a JOP that will entail looking at the dynamics in specific Helium lines. This will be submitted soon.

6. Experiment anomaly reports

Continue with these? Leave as item for SWT.

7. Publication of SOHO schedules

Purpose of SOHO planning tools is to exchange information between SOHO experiments, and to announce to the world the observing programs and targets of the SOHO experiments. SOHO Web pages exist that make the information from experiment IAPs accessible, but so far IAPs are not updated frequently enough to reflect the actual observing program, and neither do they contain information on observing targets and field-of-view (FOV).

The reason appears to be that the planning system of experimenters are not yet fully functional. B. Bentley said coordinates may be in IAPs, but not passed on to KAPs.

IAP updates: CDS keeps track of updated schedule via command uploading.

The ICD states coronal instruments only need to input coordinates, not FOV.

EIT has no planning tool, typing all in by hand. EIT planning tool cannot talk to LASCO.

UVCS is typing in things by hand. Going from sciplan, no coordinates.

Experiments and Project Scientist teams will work together to improve the situation.

8. 9. As-run database

As-run database is unclear among various teams. Luis will make it an action item to sit down with teams next week to organize this.

10. Status Data Distribution

CD-ROMs will be generated from JPL tapes with the nasty gaps removed. On Wednesday after the Daily Meeting, Luis will have a meeting about this. SWAN LOBT not always correct. Time stamp is in the data but wrong by 2 months. Real spacecraft time not on science packet. SWAN may have trouble organizing data with incorrect time stamps.

11. Special Issue: First Results from SOHO

Bernhard Fleck brought up an issue to be talked about once more at the SWT. This concerns having a joint SOHO issue in Solar Physics with early results. Discussed latest proposal of Zdenek Svestka which foresees a split of the SOHO special in 2 or 3 parts. Part I should appear in the Nov issue of Solar Physics, Part II in the January or February 97 issue. After some discussion this split solution was considered a reasonable compromise, rather than delaying the whole issue until 1997.

12. AOB

It was decided to continue having minutes of daily meetings made and distributed per subscription mailing list.

Horst Kunow suggested that an informal science workshop be held near the SWT so that the particle experiments could learn more about the science that other instruments were doing. For example, 1/2 day for a meeting then another 1/2 day of discussion. This workshop should have contributions from all scientists not just the PIs and be held around the SWT in time frame. Finally it was agreed to start with a 1/2 day workshop at the May SWT meeting (afternoon of May 16). H. Kunow volunteered to organize it.